APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday September 21, 2001 (MJD 52174) ---=== OBSERVING TEAM ===--- swing: Mike Harvanek (APO) night: Jurek Krzesinski (APO) night: Pete Newman (APO) ---=== OBSERVING PLAN ===--- Science! Spectroscopy till the moon set, then imaging. ---=== OBSERVING SUMMARY ===--- - During afternoon checkout there was a problem with id2-1 and id2-2. - MCP has been rebooted today. We verified the version v5_4_3 and use it. - during the night we closed the telescope due to the weather and reopened again latter. - note: cartridge 1 guide fiber 3 is very dirty (or perhaps some glass fibers are broken in it?). DONE two plates: - cart 1 plt 637 done with four exposures (900,900,1200,1200 sec int times). - cart 4 plt 664 done with 3 exposures (1200, 1200, 1200 sec) Two aborted: - cart 5 plt 676 - integration aborted due to the weather. S/N usless. - cart 3 plt 708 - also aborted because of the emergency closing. PROBLEMS: - TCC/MCP error no longer tracking (an old error - spontaneous abort) - happened twice. - Problems with the petal flat field screens ocuured twice during closing and reopening. (PR has been filed). - Emergency closing - again got flat field screen problems (PR has been edited and severity changed to critical high). - After a few days of silence we got again Galil full step error. ---=== OBSERVING LOG ===--- During afternoon checkout there was a problem with id2-1 and id2-2. During the test bias, both these drives shut down. See the problem section for details. MCP has been rebooted today. We verified the version v5_4_3 and use it. 19:30 - telescope open with cart 1 637 mounted. Some clouds close to the horison. 20:00 - set for the stars - we found none (previous nights offsets hasn't been cleared), but we found fk5 star and centered it with additional offests RA 5, DEC 10. Then we cleared out offsets, went to field (none stars), and found stars with a new offset RA 11, DEC 6 Note: cartridge 1 guide fiber 3 is very dirty (or perhaps some fibers are broken in it?). Seeing about 1.6'', variable. 21:00 TCC/MCP error no longer tracking (an old error - spontaneous abort), the guider failed afterwards -probably because a bad responce from TCC -see problem section #2) for details. See the murmurlog Sep 21 20:59:22 alt frame break... We aborted the exposition (10 minutes of science lost). 22:30 cart 1 plate 637 done with four exposures (total int time: 4200). Red monster present. SoS suggested postcalib, and we followed the instruction. We still see clouds, so inspite a good seeing we stick with spectroscopy. 22:40 Cart 5 plt 676 mounted. Stars found with the first attempt. 23:15 Thick clouds stopped us from observations. We've been thretened by the suddend change of the weather. 23:30 We closed the telescope, problems with the petal flat field ocuured ( petal #3?), but after pushing it down we were able to move the telescope to the stow position). 01:30 It is clearing out. 01:45 We opened the telescope with cart 4 plt 664 mounted. Again we had to bang on the petal flat field screen (it might be the petal #2, which is faulty, but we bang on the #3) to made the telescope able to move. Because of thin clouds we are using 1200 int time. 02:54 axies halted during an exposure (again), but we noticed this before the watcher did, and gotoField quickly. So we didn't abort that exposure. Seeing is getting worse: ~2.1''. 03:26 cart 4 plt 664 done with 3 exposures (total int time 3600). 03:40 cart 3 plt 708 mounted. Stars found via fk5 star. 04:05 clouds waiting. 04:20 continuing with cart 3 plt 708, focusing/centering 05:50 emergency closing!!! exposition aborted. Rain sensor went on (however no participation has been noticed, and the humidity was about 70%). Again during closing we got petal #3 error - we opened flat fields screen and then closed them, and could move the telescope. Note: after a few days of silence we've got again galil: full step error, which happened while we were closing telescope. 05:10 telescope closed. We called endNight at this point. endNight wen with no problems. ---=== IMAGING RUN SUMMARY ===--- Run Stripe Flavor Lambda Last Begin End Frame Comments ------------------------------------------------------------------ ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot --------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields nGood rowMean rowSig colMean colSig rot --------------------------------------------------------------- ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 10969, plate -9999 ------- 22:12 10969 0.0 bias {{}} 00:25 10970 0.0 bias {{}} ----- sequence 10971, plate 637 ------- 01:25 10971 0.0 bias {{}} 02:32 10972 10.0 flat pre-flat 02:34 10973 2.0 arc pre-arc 02:54 10974 900.0 target science 03:27 10975 901.0 target science science 03:49 10976 1202.0 target science Plate 637 was observed in sequence(s) 10971 for a total of 3003 seconds 04:14 10977 1201.0 target science 04:14 10977 1201.0 target science 04:14 10977 1201.0 target science 04:14 10977 1201.0 target science 04:22 10978 240.0 smear {{Smear Pattern}} 04:22 10978 240.0 smear {{Smear Pattern}} 04:22 10978 240.0 smear {{Smear Pattern}} 04:22 10978 240.0 smear {{Smear Pattern}} 04:26 10979 10.0 flat post-flat 04:26 10979 10.0 flat post-flat 04:26 10979 10.0 flat post-flat 04:26 10979 10.0 flat post-flat 04:28 10980 2.0 arc post-arc 04:28 10980 2.0 arc post-arc 04:28 10980 2.0 arc post-arc ----- sequence 10981, plate 676 ------- 04:47 10981 10.0 flat pre-flat ----- sequence 10981, plate 676 ------- 04:47 10981 10.0 flat pre-flat ----- sequence 10981, plate 676 ------- 04:47 10981 10.0 flat pre-flat 04:49 10982 2.0 arc pre-arc 04:49 10982 2.0 arc pre-arc 04:49 10982 2.0 arc pre-arc 05:17 10983 949.0 target science 05:17 10983 949.0 target science 05:17 10983 949.0 target science 05:17 10983 949.0 target science ----- sequence 10984, plate 664 ------- 08:00 10984 10.0 flat pre-flat ----- sequence 10984, plate 664 ------- 08:00 10984 10.0 flat pre-flat ----- sequence 10984, plate 664 ------- 08:00 10984 10.0 flat pre-flat 08:02 10985 2.0 arc pre-arc 08:02 10985 2.0 arc pre-arc 08:02 10985 2.0 arc pre-arc 08:34 10986 1200.0 target science 08:34 10986 1200.0 target science 08:34 10986 1200.0 target science 08:57 10987 1200.0 target science 08:57 10987 1200.0 target science 08:57 10987 1200.0 target science 09:23 10988 1201.0 target science 09:23 10988 1201.0 target science 09:23 10988 1201.0 target science 09:30 10989 240.0 smear {{Smear Pattern}} 09:30 10989 240.0 smear {{Smear Pattern}} 09:30 10989 240.0 smear {{Smear Pattern}} ----- sequence 10990, plate 708 ------- 10:04 10990 10.0 flat pre-flat ----- sequence 10990, plate 708 ------- 10:04 10990 10.0 flat pre-flat ----- sequence 10990, plate 708 ------- 10:04 10990 10.0 flat pre-flat 10:06 10991 2.0 arc pre-arc 10:06 10991 2.0 arc pre-arc 10:06 10991 2.0 arc pre-arc 10:37 10992 1200.0 target science 10:37 10992 1200.0 target science 10:37 10992 1200.0 target science ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument getclock Az Alt Rot Scale ------------------------------------------------------------------------------ 20:39 1 637 1001126379 0.004601 -0.001422 0.056273 0.999890 22:18 1 637 1001132284 0.005343 0.000913 0.054006 0.999770 22:56 1 676 1001134600 0.007177 0.001556 -0.002363 1.000090 02:26 4 664 1001147168 0.004815 0.000402 0.015295 0.999850 03:05 4 664 1001149527 0.004765 0.001185 0.013972 0.999820 04:24 3 708 1001154271 0.006225 0.000952 0.033904 1.000070 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 20:38 1 637 0.99989 1085 1964 -130 30 46.0 14.6 4 109 - 1.6 20:59 1 637 0.99989 1085 1887 -200 23 48.0 14.4 2 112 - 1.4 22:13 1 637 0.99977 2270 2779 -280 -5 50.5 14.8 1 318 - 1.2 22:52 5 676 0.99997 296 1148 -265 26 54.4 15.3 0 171 - 1.3 02:25 4 664 0.99985 1480 2285 -130 7 47.6 12.0 12 134 - 2.0 03:05 4 664 0.99982 1776 2535 -125 -8 47.6 11.0 15 138 - 2.1 04:25 3 708 1.00007 -690 490 -100 -15 56.3 11.9 12 179 - 2.1 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 18:17 63 33 1 144 (SE) 230 - 18:49 56 38 13 47 (NE) 226 - 19:21 54 38 16 67 (ENE) 485 - 19:54 54 38 11 79 (E) 227 - 20:26 57 33 3 92 (E) 245 - 20:58 57 33 1 115 (ESE) 237 - 21:28 59 24 0 77 (ENE) 209 - 21:58 59 25 0 92 (E) 229 - 22:31 59 25 0 236 (SW) 177 - 23:02 60 25 0 84 (E) 155 - 23:33 52 35 23 74 (ENE) 234 - 00:03 50 37 12 72 (ENE) 229 - 00:35 49 38 19 75 (ENE) 265 - 01:05 50 39 13 83 (E) 265 - 01:36 52 38 7 93 (E) 294 - 02:07 53 36 2 108 (ESE) 298 - 02:38 52 36 7 122 (ESE) 278 - 03:08 52 40 16 134 (SE) 743 - 03:40 51 41 12 156 (SSE) 1088 - 04:13 53 40 11 175 (S) 829 - 04:45 53 40 11 209 (SSW) 888 - 05:18 53 39 14 185 (S) 612 - ---=== TELESCOPE STATUS ===--- 19:00 telescope fidutialized, enclosure door opened, fans on. 19:30 telescope opened, cart 1 637 mounted - ready for fly. 23:30 closed due to the weather. 01:45 opened again with cart 4 664 mounted. 04:50 closed in emergency. ---=== SOFTWARE USED ===--- IOP/SOP: v3_56_9 Watcher: v2_10_0 MCP: v5_4_3 TPM: v2_3_1 AstroDa: v14_36 TCC: 2.5.5 sdssProcedures: v1_29 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES primDesOrient: 0.00 -11.80 23.00 1256.90 642.10 primOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.0 Air temp: 12.5 C Altitude: 30.00 GALILS Commanded: 1583850. 1583850. 1583850. 2600. 3800. Actual: 1583863. 1583863. 1583895. 2600. 3800. SETMIR VALUES secDesOrient: 1257.00 0.00 0.00 0.00 -175.00 secOrient: 1257.03 -0.01 0.00 0.00 -175.12 ---=== PROBLEMS IN DETAIL ===--- 1). During afternoon checkout there was a problem with id2-1 and id2-2. During the test bias, both these drives shut down. The watcher messages are given below. All drives were cleaned at the beginning of this dark run and no cleaning lights were lit on either drive. It seems a bit unlikely that both tapes were bad (although both were labelled as suspect). The instructions for recovering 2 drives were followed and id2-2 came back but id2-1 did not. The instructions for recovering a single drive were then followed and id2-1 came back. >From the watcher: dsc E archStsFail Failed to receive exabyte command completion status (dsc_E_archStsFail Failed to receive exabyte command completion status : Sep 21 16:17:27 MDT ) dsc E archShutDown1 Archiver shutting down first drive (dsc_E_archShutDown1 Archiver shutting down first drive : Sep 21 16:14:43 MDT ) dsc E trcErr dscArch CheckPrevOp: rstatus = 0x8001421c, drive 1 (dsc_E_trcErr dscArch_CheckPrevOp: rstatus = 0x8001421c, drive 1 : Sep 21 16:14:43 MDT ) archUseDrv (drive 2: archUseDrv is 0: Sep 21 16:15:43 MDT) dsc E trcErr dscArch CheckPrevOp: rstatus = 0x8001421c, drive 2 (dsc_E_trcErr dscArch_CheckPrevOp: rstatus = 0x8001421c, drive 2 : Sep 21 16:14:43 MDT ) dsc E archShutDown2 Archiver shutting down second drive (dsc_E_archShutDown2 Archiver shutting down second drive : Sep 21 16:14:43 MDT ) archEnable (archEnable is 0: Sep 21 16:15:43 MDT) dsc E archWrFail Tape write error while archiving frame 2592 6.22 (dsc_E_archWrFail Tape write error while archiving frame 2592 6.22 : Sep 21 16:14:43 MDT ) archDrvSts (drive 2: Error: Sep 21 16:21:44 MDT) 2). Spontaneous abort 0 12 I SecFocus=-200.00 0 12 I GCFocus=0.00 0 12 : Cmd='show focus' 0 12 I ScaleFac=0.999890 0 12 : Cmd='show scale' 0 12 I PrimDesOrient=1085.96, -11.80, 23.00, 1256.90, 642.10; PrimDesOrientAge= 33.2 0 12 I PrimCmdMount=-11650., -20750., -16100., -200., 31550., 30650. 0 12 I PrimActMount=-11650., -20750., -16100., -200., 31550., 30650. 0 0 F Modu=''prt_Read''; Text=''read timed out on port LTA19:'' 0 12 I PrimOrient=1085.20, -11.90, 23.25, 1256.01, 642.48 0 12 I PrimStatus=65537, 65537, 65537, 65537, 65537, 65537 0 12 I Received='' 1 lateral links (0 disengaged, 1 engaged)'' 0 12 I SecDesOrient=1886.85, 0.00, 0.00, 0.00, -259.84; SecDesOrientAge=34.8 0 12 I SecCmdMount=2377600., 2377600., 2377600., 6150., 7350. 0 12 I SecActMount=2377618., 2377436., 2377637., 6150., 7350. 0 12 I SecOrient=1886.81, -0.05, -0.05, -0.03, -260.21 0 12 I SecStatus=65537, 65537, 65537, 65537, 65537 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''AZ MOVE 23. 2840704 -0. 0058455 10799.65000''; Received=''23.352857 -2.392466 10788.094657 0 0.000000'' 0 12 I Received='' 1 piezo correction enabled (1 yes, 0 no)'' 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''ALT MOVE 47. 9940736 0. 0013917 10799.65000''; Received=''AZ MOVE 2 3.2840704 -0.0058455 10799.65000'' 0 12 I Received='' 000000470, 000001676, 000000206 piezo corrections ( microsteps)'' 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''ROT MOVE 199. 5286453 -0 .0048202 10799.65000''; Received='' '' 0 12 : Cmd='mir stat' 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''ALT STATUS''; Received='' 0048202 10799.65000'' 0 0 F Modu=''axe_o_Move''; Text=''command failed for AZ , ALT, ROT'' 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''AZ STATUS''; Received='' K'' 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''ROT STATUS''; Received='' 370 10793.992472 0 0.000000'' 0 0 I TCCStatus=''TTT'',''NNN''; TCCPos=23.35,47.98,199.58; AxePos=23.41,47.96, 199. 63 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''AZ MOVE''; Received='' ROT STATUS'' 0 0 W AzStat=NaN, NaN, NaN, NaN; AzDTime=NaN 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''ALT MOVE''; Received='' 199.555826 -1.973516 10794.117978 0 0.00000 0'' 0 0 W AltStat=NaN, NaN, NaN, NaN; AltDTime=NaN 0 0 F Modu=''prt_ReadMatch''; Text=''bad reply''; Expected=''ROT MOVE''; Received=''AZ MOVE'' 0 0 W RotStat=NaN, NaN, NaN, NaN; RotDTime=NaN; RotID=1, 0 0 0 F Modu=''axe_o_Move''; Text=''command failed for AZ , ALT, ROT'' 0 0 I SpiderInstAng = 19.611469, -0.004818, 4507844394.65000 0 0 I Modu=''axe_o_Init''; Text=''locking ports: Az , Alt, Rot'' 0 12 W AzStat=NaN, NaN, NaN, NaN; AzDTime=NaN 0 12 I AltStat=47.986801, 0.570606, 4507844394.54666, 0x00000000; AltDTime=-0.2 0 0 I Modu=''axe_o_Init''; Text=''initializing: Az , Alt, Rot'' 0 12 I RotStat=199.552730, -1.973576, 4507844394.74516, 0x00000000; RotDTime=-0 .1; RotID=1, 0 0 12 F Failed; cmdStr='axis status' 0 0 I Modu=''axe_o_Init''; Text=''setting time of: Az , Alt, Rot'' 0 0 I Modu=''axe_o_Init''; Text=''sending init files (if they exist) to: Az , Alt, Rot'' 0 0 I TCCStatus=''TTT'',''NNN''; TCCPos=23.29,47.99,199.54; AxePos=23.31,47.99, 199. 55 0 0 I SpiderInstAng = 19.611469, -0.004818, 4507844394.65000 0 0 I AzStat=23.312984, 0.000000, 4507844400.50512, 0x00000000; AzDTime=-0.1 0 0 I AltStat=47.987175, 0.000000, 4507844400.82476, 0x00000000; AltDTime=-0.1 0 0 I RotStat=199.551480, 0.000000, 4507844401.06707, 0x00000000; RotDTime=-0.1 ; RotID=1, 0 0 0 I AxisInit=''TTT'' 0 0 F Modu=''axe_i_SchMove''; Text=''tracking failed; axes halted'' 0 0 I SchMoveTimes=-4.0, -4.0, -4.0, -4.0 0 0 I ObjName='' '' 0 0 F Modu=''exe_Track''; Text=''no longer tracking''